# Current Status of Inert Higgs Dark Matter with Dark Fermions

### v1

@@ II The model
Here $D_{\mu}$ is the covariant derivative in the SM, and the $SU(2)_{L}$ contraction $\Psi\cdot H_{\alpha}=\epsilon_{ab}\Psi_{a}H_{\alpha b}$ with $\epsilon_{ab}=(i\sigma_{2})_{ab}$ and $\alpha=1,2$. Note that $M_{\Psi}$, $M_{\chi}$, $\mu_{i}$ ($i=1,2$) are parameters with the same dimension as mass and $y_{H_{1}}$, $\lambda_{i}$ ($i=1-5$), $\kappa_{M}$, $\kappa_{\mu}$ are dimensionless parameters. Furthermore, we assume all dimensionless parameters are real in this work.
@@
  • The singlet fermion field $\chi$ cannot be a Majorana type as it would lead to excessively large neutrino masses via the one-loop process Ma:2006km in our parameter space of interest.
@@
  • We only consider the second generation of the lepton sector in ${\cal L}_{\text{mix}}$ for simplicity, but it can be straightforwardly extended to include all three generations of the lepton sector.
@@
Here we impose the stipulation that $S$ as the DM candidate, ensuring that $m_{S}<m_{n_{1},n_{2}}$. Therefore, $m_{n_{1}}>100$ GeV is required in our scanning. The selection of the scanning ranges for $m_{S}$, $\Delta^{0}$, $\Delta^{\pm}$, and $\lambda_{S}$ adheres to the guidelines outlined in our previous investigations Fan:2022dck ; Tsai:2019eqi . Additionally, for perturbative consistency, we enforce the bounds $\lvert y_{H_{1}}\rvert$, $\lvert\kappa_{M}\rvert$, and $\lvert\kappa_{\mu}\rvert$ to be less than $\sqrt{4\pi}$. Furthermore, to evade the constraints imposed by ATLAS and CMS slepton searches ATLAS:2019lff ; CMS:2020bfa , we conservatively require the condition $M_{\Psi}>600$ GeV.
III The extra $(g-2)_{\mu}$, $M_{W}$ contributions and other constraints
In this section, we will initially delve into the manner in which our model contributes to $\Delta a_{\mu}$ and $M_{W}$, as outlined in Sec. III.1. Subsequently, we will provide an overview of the relevant theoretical and experimental constraints in Sec. III.2.
III.1 Contributions to $\Delta a_{\mu}$ and $M_{W}$
@@
Figure 2: The contribution to $\Delta a_{\mu}$ in this model on the $(m_{\Psi},m_{H^{\pm}})$ plane. We fixed $\Delta M_{\Psi\chi}=100$ GeV, $|y_{H_{1}}|=1$ and $|\kappa_{M}\times\kappa_{\mu}|=0.5$. The sign of $y_{H_{1}}\kappa_{M}\kappa_{\mu}$ is fixed to be negative in the left panel while it is positive in the right panel. The solid blue, green, dotted blue and green represent the values of $\Delta a_{\mu}$ of $5\times 10^{-9}$, $10^{-9}$, $-10^{-9}$ and $-5\times 10^{-9}$, respectively. The dash-dotted black line indicates $\Delta a_{\mu}=0$, where a complete cancellation between the two terms in the brackets of Eq. 9 occurs. The orange band is the $1\sigma$ region favored to the discrepancy between the current experiment measurements and the SM results where the HVP is calculated using the data-driven method Aoyama:2020ynm . The light blue region is excluded by the LHC constraints.
@@
Figure 4: Contribution to $\Delta m_{W}$ from the $\Psi$ field projected onto the $(m_{\Psi},|y_{H_{1}}|)$ plane. In the left panel, we fixed $\Delta M_{\Psi\chi}=1$ GeV, while in the right panel, $\Delta M_{\Psi\chi}=200$ GeV. The dotted red, dotted blue, solid blue and solid green lines represent the values of $\Delta m_{W}$ equal to $-50$ MeV, $-25$ MeV, $25$ MeV, and $75$ MeV, respectively. The dash-dotted black line indicates $\Delta m_{W}=0.0$.
@@
${\cal S}$ $=$ $\frac{g^{2}s^{2}_{W}c^{2}_{W}}{4\pi^{2}\alpha_{\text{EM}}}\left[\frac{1}{3}\left\{\ln\left(\frac{\Lambda_{\text{EW}}^{2}}{m_{\psi^{\pm}}^{2}}\right)-\cos^{4}\alpha\ln\left(\frac{\Lambda_{\text{EW}}^{2}}{m_{n_{2}}^{2}}\right)-\sin^{4}\alpha\ln\left(\frac{\Lambda_{\text{EW}}^{2}}{m_{n_{1}}^{2}}\right)\right\}-\frac{1}{2}\sin^{2}2\alpha\left\{\ln\left(\frac{\Lambda_{\text{EW}}^{2}}{m_{n_{1}}m_{n_{2}}}\right)\right.\right.$ (14)
$\left.\left.+\frac{m_{n_{1}}^{4}-8m_{n_{1}}^{2}m_{n_{2}}^{2}+m_{n_{2}}^{4}}{9(m_{n_{1}}^{2}-m_{n_{2}}^{2})^{2}}+\frac{(m_{n_{1}}^{2}+m_{n_{2}}^{2})(m_{n_{1}}^{4}-4m_{n_{1}}^{2}m_{n_{2}}^{2}+m_{n_{2}}^{4})}{6(m_{n_{1}}^{2}-m_{n_{2}}^{2})^{3}}\ln\left(\frac{m_{n_{2}}^{2}}{m_{n_{1}}^{2}}\right)\right.\right.$
@@
III.2 Theoretical and experimental constraints
  • Production of inert scalars at the LEP
  • Higgs boson invisible and exotic decays
    When the inert scalars have masses lower than half of $m_{h}$, certain decay modes become kinematically possible. These include the Higgs invisible decay, $h\rightarrow SS$, and the Higgs exotic decays, $h\rightarrow AA,H^{+}H^{-}$ [^1] Given that the soft visible objects cannot be identified at the LHC, we classify $h\rightarrow AA,H^{+}H^{-}$ as Higgs invisible decays when $\Delta^{0,\pm}$ is less than about $2$ GeV ATLAS:2019lng .. In our analysis, we consider BR$(h\rightarrow\text{invisible})<10.7\%$ ATLAS:2023tkt and BR$(h\rightarrow\text{undetected})<19\%$ ATLAS:2020qdt at $95\%$ C.L.
  • The $h\rightarrow\gamma\gamma$ signal strength
    $\mu_{\gamma\gamma}=1.04^{+0.10}_{-0.09}(1.12\pm 0.09).$ (17)
  • Mono-jet and compressed mass spectra searches at the LHC
  • DM relic density
    $\Omega_{\text{DM}}h^{2}=0.1198\pm 0.0015.$ (18)
  • DM direct detection
@@ IV.1 Methodology
where $\chi^{2}_{\mathrm{a_{\mu}}}$, $\chi^{2}_{\mathrm{m_{w}}}$, $\chi^{2}_{\mathrm{\Omega h^{2}}}$, $\chi^{2}_{\mathrm{h_{exo}}}$, $\chi^{2}_{\mathrm{h_{inv}}}$, $\chi^{2}_{\mathrm{R_{\gamma\gamma}}}$ and $\chi^{2}_{\mathrm{LZ}}$ represent the $\chi^{2}$ from muon $g-2$, $W$-boson mass, DM relic density, Higgs boson exotic decays, Higgs boson invisible decays, signal strength of Higgs boson to diphoton and DM direct detection, respectively.
To elucidate the effects of quality anomalies from muon $g-2$ and $W$-boson mass, we conduct four sets of numerical scans:
@@
Here $m_{W,\text{CDF II}}$ and $m_{W,\text{PDG}}$ are the $W$ boson mass results from the CDF II experiments and from the Particle Data Group ParticleDataGroup:2022pth , respectively. The muon $g-2$ results, $\Delta a_{\mu}^{e^{+}e^{-}\text{ HVP}}$ and $\Delta a_{\mu}^{\text{Lattice HVP}}$, correspond to the HVP contributions obtained via data-driven and lattice calculation methods, respectively.
On the other hand, we also compare our allowed parameter space with several DM indirect detection constraints. These include the bounds established by the Fermi LAT observations of dwarf Spheroidal galaxies Fermi-LAT:2016uux , the signal regions of the Fermi LAT Galactic Center $\gamma$-ray excess Hooper:2010mq ; Zhou:2014lva ; Calore:2014xka ; Daylan:2014rsa , and the Alpha Magnetic Spectrometer AMS-02 experiment antiproton excess 1610.03840 ; 1610.03071 ; 1803.02163 ; 1903.02549 .
@@ IV.2 Results
[…]
In the left panel of Fig. 6, we present the 95% C.L. regions for the Higgs-DM-DM coupling $\lambda_{S}$ as a function of the DM mass. We find that $\lambda_{S}$ must remain small, mainly due to the Planck relic density constraint, as a larger coupling leads to an overabundance of DM. As expected, the viable range of $\lambda_{S}$ is not significantly affected when the $m_{W,\text{PDG}}+\Delta a_{\mu}^{\text{Lattice HVP}}$ scenario is imposed, as shown by the gray region. Notably, around $m_{S}\simeq 62$ GeV, the coupling $\lambda_{S}$ exhibits a bottleneck shape, a feature driven by the Higgs resonance mechanism. As the DM mass moves away from the resonance, increasingly larger values of $\lambda_{S}$ are needed to open up other annihilation channels.
[…]
Figure 7: Left panel: The $95\%$ C.L. regions from all new constraints projected on the ($m_{S}$, $\sigma_{p}^{\text{SI}}$) panel. The solid black and brown lines represent the current constraints set by PandaX-4T (2021) PandaX-4T:2021bab and LZ 2022 LZ:2022ufs , respectively. The dashed blue and green lines are the expected detection sensitivity of XENONnT/LZ XENON:2020kmp ; 1509.02910 and DARWIN Schumann:2015cpa . Right panel: The $95\%$ C.L. regions on the ($m_{S}$, $\langle\sigma v\rangle$) panel. The dotted green and dashed blue contours indicate the parameter regions that can explain the GCE and AMS-02 $\overline{p}$ excesses through the DM annihilation channel $SS\to WW^{*}$ Zhu:2022tpr .
@@
Figure 8: Left panel: The total production cross section, including $pp\to AS$, $pp\to AH^{\pm}$, $pp\to SH^{\pm}$, and $pp\to H^{\pm}H^{\mp}$, is calculated with DM mass $m_{S}=60$ GeV while varying the mass splitting $\Delta^{0,\pm}$ at $\sqrt{s}=13$ TeV. Right panel: The production cross section of $pp\to n_{1}\overline{n_{1}}$ at $\sqrt{s}=13$ TeV and $\mu^{+}\mu^{-}\to n_{1}\overline{n_{1}}$ at $\sqrt{s}=3$ TeV with varying the mass of $n_{1}$.
@@
    Decoupled $n_{1}$: In this scenario, the heavier neutral scalar $A$ produced at the LHC can decay into a pair of leptons, with the invariant mass of these leptons being sensitive to the mass splitting $\Delta^{0}$. We recast the current exclusion limit from the CMS SUSY compressed mass spectrum searches CMS:2021edw via MadAnalysis5. Specifically, we focus on the exclusive signal region, $2l$-Ewk SR with $p^{\text{miss}}_{T}=200-240$ [GeV] and $M(ll)=1-4$ [GeV] as outlined in Table 4 of Ref. CMS:2021edw . In this region, $2$ events were observed, while $5.5\pm 2.5$ total background events were expected. This exclusive signal region provides the most stringent constraints. We consider all possible signal processes that produce soft leptons signatures, including $SA$, $AH^{\pm}$, and $H^{\pm}H^{\mp}$. As shown in the left panel of Fig. 8, the combined cross sections for these processes, without the decays of $A$ and $H^{\pm}$, are on the order of a few pb. Moreover, the decay branching ratios of $H^{\pm}$ are approximately $BR(H^{\pm}\to Sjj^{\prime})\sim 0.67$ and $BR(H^{\pm}\to Sl^{\pm}\nu)\sim 0.33$. Similarly, the branching ratios for $A$ are $BR(A\to Sjj)\sim 0.66$, $BR(A\to Sl^{+}l^{-})\sim 0.12$, and $BR(A\to S\nu\overline{\nu})\sim 0.22$. In the parameter space of interest, these branching ratios show slight dependence on the mass splitting $\Delta^{0,\pm}$. However, these cross sections are still not large enough to exclude the parameter space for the dominant $SA$ co-annihilaiton region, except for $m_{S}\lesssim 60$ GeV with $\Delta^{0}<10$ GeV, as indicated by the black solid line in the upper-right panel of Fig. 5. Note that $\Delta^{\pm}=50$ GeV is fixed in this analysis. The magnitude of the mass splitting inversely correlates with the stringency of the imposed constraints. Specifically, larger mass splittings result in less stringent restrictions, while smaller mass splittings enforce stricter constraints. However, the overall impact on the exclusion line remains relatively modest. As we will see in Sec. IV.3, the upcoming luminosity upgrade of the LHC will thoroughly test the entire $SA$ co-annihilation region.

### v2

@@ II The model
Here $D_{\mu}$ is the covariant derivative in the SM, and the $SU(2)_{L}$ contraction $\Psi\cdot H_{\alpha}=\epsilon_{ab}\Psi_{a}H_{\alpha b}$ with $\epsilon_{ab}=(i\sigma_{2})_{ab}$ and $\alpha=1,2$. Note that $M_{\Psi}$, $M_{\chi}$, $\mu_{i}$ ($i=1,2$) are parameters with the same dimension as mass and $y_{H_{1}}$, $\lambda_{i}$ (i=15i=1-5), $\kappa_{M}$, $\kappa_{\mu}$ are dimensionless parameters. Furthermore, we assume all dimensionless parameters are real in this work [^1] Note that the Yukawa-type couplings $y_{H_{1}}$, $\kappa_{M}$, and $\kappa_{\mu}$ in Eqs. 2 and 4 are generally complex..
@@
  • The singlet fermion field $\chi$ in this model can be either Dirac-type or Majorana-type. However, if $\chi$ is Majorana-type, as discussed in Ref. Ma:2006km ; Cai:2017jrq , the parameter space under consideration would result in excessively large neutrino masses via the one-loop process. Thus, from a phenomenological perspective, the $(g-2)_{\mu}$ excess provides a strong indication for treating the $\chi$ field as Dirac-type rather than Majorana-type in this study.
@@
  • In the minimal model, ${\cal L}_{\text{mix}}$, we consider only the second generation of the lepton sector. However, it can be straightforwardly extended to include all three generations. The couplings of the dark fermion to $e$, $\mu$, and $\tau$ can be treated as free and independent parameters. This flexibility allows us to set the couplings to $e$ and $\tau$ small enough that they do not affect the results of this analysis.
@@
Here we impose the stipulation that $S$ as the DM candidate, ensuring that $m_{S}<m_{n_{1},n_{2}}$. In some specific parameter space, the neutral particle $n_{1}$ can be regarded as the next-to-lightest dark particle, and its possible mass range is discussed in Sec. IV.2. The selection of the scanning ranges for $m_{S}$, $\Delta^{0}$, $\Delta^{\pm}$, and $\lambda_{S}$ adheres to the guidelines outlined in our previous investigations Fan:2022dck ; Tsai:2019eqi . Additionally, for perturbative consistency, we enforce the bounds $\lvert y_{H_{1}}\rvert$, $\lvert\kappa_{M}\rvert$, and $\lvert\kappa_{\mu}\rvert$ to be less than $\sqrt{4\pi}$. Furthermore, to evade the constraints imposed by ATLAS and CMS slepton searches ATLAS:2019lff ; CMS:2020bfa , we conservatively require the condition $M_{\Psi}>600$ GeV.
III Other constraints and extra $(g-2)_{\mu}$, $M_{W}$ contributions
In this section, we will provide an overview of the relevant theoretical and experimental constraints in Sec. III.1. Subsequently, we will initially delve into the manner in which our model contributes to $\Delta a_{\mu}$ and $M_{W}$, as outlined in Sec. III.2.
III.1 Theoretical and experimental constraints
  • Production of inert scalars at the LEP
  • Higgs boson invisible and exotic decays
[…]
  • The $h\rightarrow\gamma\gamma$ signal strength
    $\mu_{\gamma\gamma}=1.04^{+0.10}_{-0.09}(1.12\pm 0.09).$ (9)
  • Mono-jet and compressed mass spectra searches at the LHC
  • DM relic density
    $\Omega_{\text{DM}}h^{2}=0.1198\pm 0.0015.$ (10)
  • DM direct detection
III.2 Contributions to $\Delta a_{\mu}$ and $M_{W}$
@@
Figure 2: The contribution to $\Delta a_{\mu}$ in this model on the $(m_{\Psi},m_{H^{\pm}})$ plane. We fixed $\Delta M_{\Psi\chi}=100$ GeV, $|y_{H_{1}}|=1$ and $|\kappa_{M}\times\kappa_{\mu}|=0.5$. The sign of $y_{H_{1}}\kappa_{M}\kappa_{\mu}$ is fixed to be negative in the left panel while it is positive in the right panel. The solid blue, green, dotted blue and green represent the values of $\Delta a_{\mu}$ of $5\times 10^{-9}$, $10^{-9}$, $10^{-9}$ and $5\times 10^{-9}$, respectively. The dash-dotted black line indicates $\Delta a_{\mu}=0$, where a complete cancellation between the two terms in the brackets of Eq. 11 occurs. The orange band is the $1\sigma$ region favored to the discrepancy between the current experiment measurements and the SM results where the HVP is calculated using the data-driven method Aoyama:2020ynm . The light blue region is excluded by the LHC constraints.
@@
Figure 4: Contribution to $\Delta m_{W}$ from the $\Psi$ field projected onto the $(m_{\Psi},|y_{H_{1}}|)$ plane. In the left panel, we fixed $\Delta M_{\Psi\chi}=1$ GeV, while in the right panel, $\Delta M_{\Psi\chi}=200$ GeV. The dotted red, dotted blue, solid blue and solid green lines represent the values of $\Delta m_{W}$ equal to 50-50 MeV, 25-25 MeV, $25$ MeV, and $75$ MeV, respectively. The dash-dotted black line indicates $\Delta m_{W}=0.0$.
@@
[…]
@@ IV.1 Methodology
where $\chi^{2}_{\mathrm{a_{\mu}}}$, $\chi^{2}_{\mathrm{m_{w}}}$, $\chi^{2}_{\mathrm{\Omega h^{2}}}$, $\chi^{2}_{\mathrm{h_{exo}}}$, $\chi^{2}_{\mathrm{h_{inv}}}$, $\chi^{2}_{\mathrm{R_{\gamma\gamma}}}$ and $\chi^{2}_{\mathrm{LZ}}$ represent the $\chi^{2}$ from muon g2g-2, $W$-boson mass, DM relic density, Higgs boson exotic decays, Higgs boson invisible decays, signal strength of Higgs boson to diphoton and DM direct detection, respectively.
To elucidate the effects of quality anomalies from muon g2g-2 and $W$-boson mass, we conduct four sets of numerical scans:
@@
Here $m_{W,\text{CDF II}}$ and $m_{W,\text{PDG}}$ are the $W$ boson mass results from the CDF II experiments and from the Particle Data Group ParticleDataGroup:2022pth , respectively. The muon g2g-2 results, $\Delta a_{\mu}^{e^{+}e^{-}\text{ HVP}}$ and $\Delta a_{\mu}^{\text{Lattice HVP}}$, correspond to the HVP contributions obtained via data-driven and lattice calculation methods, respectively.
On the other hand, we also compare our allowed parameter space with several DM indirect detection constraints. These include the bounds established by the Fermi LAT observations of dwarf Spheroidal galaxies Fermi-LAT:2016uux , the signal regions of the Fermi LAT Galactic Center $\gamma$-ray excess Hooper:2010mq ; Zhou:2014lva ; Calore:2014xka ; Daylan:2014rsa , and the Alpha Magnetic Spectrometer AMS-02 experiment antiproton excess Cui:2016ppb ; Cuoco:2016eej ; Cui:2018klo ; Cholis:2019ejx .
@@ IV.2 Results
Within the range of model parameters we consider, the DM mass $m_{S}$ is found to lie between 54 and $74\,\text{GeV}$, irrespective of whether the anomalies in the muon magnetic moment and the $W$ boson mass are taken into account. In the upper left panel of Fig. 5, we show the $95\%$ C.L. region for $m_{S}$ and $\Delta^{0}$. The DM mass clearly falls into three distinct regions, each corresponding to a different dominant DM annihilation mechanism in the early universe. For $m_{S}\approx m_{h}/2$, this corresponds to the Higgs resonance. The $\Delta^{0}\lesssim 10\,\text{GeV}$ indicates $SA$ co-annihilation, while $71\lesssim m_{S}\lesssim 74\,\text{GeV}$ corresponds to the $SS\to WW^{*}$ process. The gray region contains a few scattered points, which arise from DM annihilation processes dominated by neutrino final states via the exchange of $n_{1,2}$. However, the significant overlap between the gray and red regions suggests that the anomalies in the muon magnetic moment and the $W$ boson mass have only a slight effect on the dominant DM annihilation mechanisms.
For the $SA$ co-annihilation mechanism, we present an enlarged view of the $95\%$ C.L. regions for $m_{S}$ versus $\Delta^{0}$ in the upper right panel of Fig. 5. The solid black line represents the exclusion limits from the CMS search for two or three soft leptons at the $13\,\text{TeV}$ LHC, with an integrated luminosity of $137\,\text{fb}^{-1}$ CMS:2021edw . The green dashed line illustrates the projected $2\sigma$ sensitivity, assuming a conservative background uncertainty of $\sigma_{b}/b=45\%$, for an integrated luminosity of $300\,\text{fb}^{-1}$ in one of the signal regions. With the forthcoming LHC luminosity upgrades, the entire $SA$ coannihilation region is expected to be thoroughly tested. Detailed calculations and discussions regarding collider experiments can be found in the latter part of this section and in Sec. IV.3.
In the left panel of Fig. 6, we present the 95% C.L. regions for the Higgs-DM-DM coupling $\lambda_{S}$ as a function of the DM mass. We find that $\lambda_{S}$ must remain small, mainly due to the Planck relic density constraint, as a larger coupling leads to a shortage in DM abundance. As expected, the viable range of $\lambda_{S}$ is not significantly affected when the $m_{W,\text{PDG}}+\Delta a_{\mu}^{\text{Lattice HVP}}$ scenario is imposed, as shown by the gray region. Notably, around $m_{S}\simeq 62$ GeV, the coupling $\lambda_{S}$ exhibits a bottleneck shape, a feature driven by the Higgs resonance mechanism. As the DM mass moves away from the resonance, increasingly larger values of $\lambda_{S}$ are needed to open up other annihilation channels.
Figure 7: The $95\%$ C.L. regions for the neutral dark particle mass $m_{n_{1}}$ and the DM mass $m_{S}$. The color definitions are consistent with those in Fig. 6.
@@
Figure 8: Left panel: The $95\%$ C.L. regions from all new constraints projected on the ($m_{S}$, $\sigma_{p}^{\text{SI}}$) panel. The solid black and brown lines represent the current constraints set by PandaX-4T (2021) PandaX-4T:2021bab and LZ 2022 LZ:2022ufs , respectively. The dashed blue and green lines are the expected detection sensitivity of XENONnT/LZ XENON:2020kmp ; LZ:2015kxe and DARWIN Schumann:2015cpa . Right panel: The $95\%$ C.L. regions on the ($m_{S}$, $\langle\sigma v\rangle$) panel. The dotted green and dashed blue contours indicate the parameter regions that can explain the GCE and AMS-02 $\overline{p}$ excesses through the DM annihilation channel $SS\to WW^{*}$ Zhu:2022tpr .
@@
Figure 9: Left panel: The total production cross section, including $p\to AS, pp$$\to AH^{\pm}, pp$$p\to SH^{\pm}$, and $pp\to H^{\pm}H^{\mp}$, is calculated with DM mass $m_{S}=60$ GeV while varying the mass splitting $\Delta^{0,\pm}$ at $\sqrt{s}=13$ TeV. Right panel: The production cross section of $pp\to n_{1}\overline{n_{1}}$ at $\sqrt{s}=13$ TeV and $\mu^{+}\mu^{-}\to n_{1}\overline{n_{1}}$ at $\sqrt{s}=3$ TeV with varying the mass of $n_{1}$.
@@
    Decoupled $n_{1}$: In this scenario, the heavier neutral scalar $A$ produced at the LHC can decay into a pair of leptons, with the invariant mass of these leptons being sensitive to the mass splitting $\Delta^{0}$. We recast the current exclusion limit from the CMS SUSY compressed mass spectrum searches CMS:2021edw via MadAnalysis5. Specifically, we focus on the exclusive signal region, $2l$-Ewk SR with $p^{\text{miss}}_{T}=200-240$ [GeV] and M(ll)=14M(ll)=1-4 [GeV] as outlined in Table 4 of Ref. CMS:2021edw . In this region, $2$ events were observed, while $5.5\pm 2.5$ total background events were expected. This exclusive signal region provides the most stringent constraints. We consider all possible signal processes that produce soft leptons signatures, including $SA$, $AH^{\pm}$, and $H^{\pm}H^{\mp}$. As shown in the left panel of Fig. 9, the combined cross sections for these processes, without the decays of $A$ and $H^{\pm}$, are on the order of a few pb. Moreover, the decay branching ratios of $H^{\pm}$ are approximately $BR(H^{\pm}\to Sjj^{\prime})\sim 0.67$ and $BR(H^{\pm}\to Sl^{\pm}\nu)\sim 0.33$. Similarly, the branching ratios for $A$ are $BR(A\to Sjj)\sim 0.66$, $BR(A\to Sl^{+}l^{-})\sim 0.12$, and $BR(A\to S\nu\overline{\nu})\sim 0.22$. In the parameter space of interest, these branching ratios show slight dependence on the mass splitting $\Delta^{0,\pm}$. However, these cross sections are still not large enough to exclude the parameter space for the dominant $SA$ co-annihilaiton region, except for $m_{S}\lesssim 60$ GeV with $\Delta^{0}<10$ GeV, as indicated by the black solid line in the upper-right panel of Fig. 5. Note that $\Delta^{\pm}=50$ GeV is fixed in this analysis [^3] We observe that the signal efficiency slightly increases for enhanced $\Delta^{\pm}$ when $m_{S}\gtrsim 70$ GeV. However, these effects are minimal, and no allowed parameter space exists in those regions.. The magnitude of the mass splitting inversely correlates with the stringency of the imposed constraints. Specifically, larger mass splittings result in less stringent restrictions, while smaller mass splittings enforce stricter constraints. However, the overall impact on the exclusion line remains relatively modest. As we will see in Sec. IV.3, the upcoming luminosity upgrade of the LHC will thoroughly test the entire $SA$ co-annihilation region.